# Error Propagation Mechanisms and Compensation Strategies for Quantized Diffusion Models

### v1

@@ 3.3 Timestep-Aware Compensation
In this section, we address Issue 2 to derive the definitive form of the correction terms. By visualizing and analyzing the noise estimation of the full-precision model ($\mu_{\theta}(\tilde{x}_{t},t)$), the noise estimation of the quantized model ($\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$), and the output distortion ($\varepsilon_{t}$) across different time steps, two critical empirical observations emerge:Timestep-Dependent Error Characteristics—the per-step quantization error $\varepsilon_{t}$ exhibits significant variations across timesteps, with distinct spatial and magnitude patterns at different stages of the denoising process. Output-Correlated Error Propagation—a strong statistical correlation exists between $\varepsilon_{t}$ and the quantized model’s outputs $\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$, particularly in high-frequency regions.
@@
where $\mathbf{K}_{t}\in\mathbb{R}^{C}$ denotes a timestep-conditioned channel-wise scaling matrix, and $\odot$ represents element-wise multiplication. We now focus on finding a loss function $L$, by minimizing which, we can efficiently reconstruct $\mu_{\theta}(\tilde{x}_{t},t)$, $\forall t\in[0,T]$. We adopt the Mean Squared Error (MSE) as the loss function and introduce regularization terms to prevent overfitting
@@
Based on the provided calibration data, we can cache the noise prediction outputs of the full-precision model and the quantized model, and then compute $K\in\mathbb{R}^{T\times C}$ offline, which is directly used for quantization errors reconstruction during inference. Different values of $\lambda_{1}$ correspond to different values of $K$. We offer two methods, grid search and empirical rule, to determine the optimal value and compare their advantages and disadvantages. Finally, the latter is adopted, that means $\lambda_{1}=0.01\times\frac{\text{mean}(\tilde{\mu}^{2})}{\text{var}(\mu)}$. Substituting Eq.14 into Eq.13, we obtain the final form of the closed-form solution for cumulative error:
@@ 4.1 Implementation Details
Quantization Scheme. SVDQuantsvdqunat introduces an additional low-rank branch that can mitigate quantization challenges in both weights and activations, establishing itself as a new benchmark for PTQ algorithms. In this study, we build our quantization strategy upon it by integrating cumulative error correction mechanisms. This approach ensures that performance comparisons remain unaffected by operator-level quantization configurations. In the 8-bit configuration, our approach employs per-token dynamic quantization for activations and per-channel weight quantization, complemented by a low-rank auxiliary branch with a rank of 16. For the 4-bit configuration, we apply per-group symmetric quantization to both activations and weights, using a low-rank branch with rank 32 and setting the group size to 64. All nonlinear activation and normalization layers are not quantized, meanwhile the inputs of linear layers in adaptive normalization are kept in 16 bits. To comprehensively evaluate the effectiveness of TCEC, we conducted comparative experiments with recent SOTA quantization algorithms across diverse generation tasks.
Image Genration Evaluation. We benchmark TCEC using SDXLpodell2023sdxl , SDXL-Turbopodell2023sdxl and PixArt modelschen2023pixart chen2024pixart including both the UNet and DiT backbones. SDXL-Turbo uses the default configuration of 4 inference steps, while SDXL employs the DDIM sampler with 50 steps.Since PixArt utilizes the DPM++ solver, we adapted TCEC to this advanced high-order solver to demonstrate its compatibility across different solvers. The detailed validation process is presented in Appendix 7.2. To precompute the channel-wise scaling matrices $K$ and $\lambda_{1}$ mentioned in section3.3, we randomly sampled 1,024 prompts from COCOchen2015microsoft as the calibration dataset. To evaluate the generalization capability of TCEC, we sample 5K prompts from the MJHQ-30Kli2024playground and the summarized Densely Captioned Images(sDCI)urbanek2024picture for benchmarking.
Video Genration Evaluation. We apply TCEC to OpenSORAzheng2024open , the videos are generated with 100-steps DDIM with CFG scale of 4.0. We evaluate the quantized model on VBenchhuang2024vbench to provide comprehensive results. Following prior researchren2024consisti2v , we evaluate video quality from three distinct dimensions using eight carefully selected metrics. Aesthetic Quality and Imaging Quality focus on assessing the quality of individual frames, independent of temporal factors. Subject Consistency, Background Consistency, Motion Smoothness, and Dynamic Degree measure cross-frame temporal coherence and dynamics. Finally, Scene Consistency and Overall Consistency gauge the alignment of the video with the user-provided text prompt. We collected 128 samples from WebVidnan2024openvid as the calibration dataset to calculate $K$ and $\lambda_{1}$.
@@ 4.2 Main Results
Image Genration Evaluation. As shown in Table 1, we conduct extensive experiments at two quantization precisions: W8A8 and W4A4. We observe that, across all precision levels, TCEC achieves better image fidelity and diversity, and it even matches the 16-bit results under W8A8 quantization. For UNet-based models, on SDXL, our W4A4 model substantially outperforms SVDQuant W4A4, the current SOTA 4-bit approach, achieving an sDCI PSNR of 21.9. This even surpasses TensorRT’s W8A8 result of 21.7, demonstrating robust performance under lower-bit quantization. On SDXL-Turbo, MixDQ W4A4 exhibits abnormal FID and IR metrics, indicating quantization failure. This highlights the greater difficulty of quantizing models with a small number of inference steps. Our W4A4 model surpasses SVDQuant by 0.017 and 0.02 in MJHQ IR and sDCI IR metrics, respectively, suggesting a stronger alignment with human preferences. The larger performance gains of TCEC on SDXL than SDXL-Turbo highlight the critical role of cumulative error correction in longer inference sequences. For DiT-based model, on PixArt-$\sum$, our W4A4 model significantly surpasses SVDQuant’s W4A4 results across all metrics. Leveraging the DPM++ solver, PixArt-$\sum$ demonstrates TCEC’s robustness across different solver configurations. As shown in Figure 2, when compared with ViDiT-Q W4A4 and SVDQuant W4A4, the TECE W4A4 method demonstrates less quality degradation and smaller changes in image content. More visual results can be found in Appendix D.
@@
Video Genration Evaluation. As shown in Table 2, our evaluation compares SOTA PTQ methods under both W8A8 and W4A8 configurations. In Imaging Quality, TCEC achieves 65.56 (W8A8) and 64.97 (W4A8), surpassing ViDiT-Q by +2.08 and +3.90 absolute points respectively. It demonstrates significant advantages in frame-wise quality metrics that evaluate static visual fidelity independent of temporal factors. TCEC achieves 96.27 motion smoothness at 8-bit and 96.01 at 4-bit quantization, outperforming ViDiT-Q by 0.13 and 0.32 respectively. This validates the effectiveness of our temporal-channel decoupled compensation strategy in handling error accumulation across denoising steps. The scene consistency metric reaches 39.58 (8-bit) and 37.20 (4-bit), establishing 1.36 and 1.01 improvements over baselines, which confirms stable long-sequence generation through temporal error propagation modeling.
@@
Hardware Resource Savings. TCEC constructs its quantization strategy based on SVDQuant by integrating cumulative error correction mechanisms. As TCEC’s research focus is not on engineering performance, we fully reuse its inference engine, Nunchakusvdqunat . The cumulative error correction terms introduce only two sets of latent transformations, and experimental tests show that the resulting end-to-end (E2E) latency degradation is less than $1\%$. Therefore, on one 12B model, TCEC-W4A4 reduces the memory footprint by $3.5\times$ compared to the FP16 model and accelerates inference by $3\times$ versus NF4 weight-only quantization on the laptop-level RTX 4090, with engineering performance comparable to SVDQuant and significantly outperforming other PTQ algorithms.
@@
Method Bit-width W/A Imaging Quality Aesthetic Quality Motion Smooth. Dynamic Degree BG. Consist. Subject Consist. Scene Consist. Overall Consist. - 16/16 63.68 57.12 96.28 56.94 96.13 90.28 39.61 26.21 Q-Diffusion 8/8 60.38 55.15 94.44 68.05 94.17 87.74 36.62 25.66 Q-DiT 8/8 60.35 55.80 93.64 68.05 94.70 86.94 32.34 26.09 PTQ4DiT 8/8 56.88 55.53 95.89 63.88 96.02 91.26 34.52 25.32 SmoothQuant 8/8 62.22 55.90 95.96 68.05 94.17 87.71 36.66 25.66 Quarot 8/8 60.14 53.21 94.98 66.21 95.03 85.35 35.65 25.43 ViDiT-Q 8/8 63.48 56.95 96.14 61.11 95.84 90.24 38.22 26.06 TCEC 8/8 65.56 57.12 96.27 61.09 96.23 91.34 39.58 26.20 Q-DiT 4/8 23.30 29.61 97.89 4.166 97.02 91.51 0.00 4.985 PTQ4DiT 4/8 37.97 31.15 92.56 9.722 98.18 93.59 3.561 11.46 SmoothQuant 4/8 46.98 44.38 94.59 21.67 94.36 82.79 26.41 18.25 Quarot 4/8 44.25 43.78 92.57 66.21 94.25 84.55 28.43 18.43 ViDiT-Q 4/8 61.07 55.37 95.69 58.33 95.23 88.72 36.19 25.94 TCEC 4/8 64.97 56.90 96.01 59.42 97.01 90.05 37.20 26.20

### v2

@@ 3.3 Timestep-Aware Compensation
In this section, we address Issue 2 to derive the definitive form of the correction terms. By visualizing and analyzing the noise estimation of the full-precision model ($\mu_{\theta}(\tilde{x}_{t},t)$), the noise estimation of the quantized model ($\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$), and the output distortion ($\varepsilon_{t}$) across different time steps in Appendix E, two critical empirical observations emerge:Timestep-Dependent Error Characteristics—the per-step quantization error $\varepsilon_{t}$ exhibits significant variations across timesteps, with distinct spatial and magnitude patterns at different stages of the denoising process. Output-Correlated Error Propagation—a strong statistical correlation exists between $\varepsilon_{t}$ and the quantized model’s outputs $\tilde{\mu}_{\theta}(\tilde{x}_{t},t)$, particularly in high-frequency regions.
@@
where $\mathbf{K}_{t}\in\mathbb{R}^{C}$ denotes a timestep-conditioned channel-wise scaling matrix, and $\odot$ represents element-wise multiplication. We now focus on finding a loss function $L$, by minimizing which, we can efficiently reconstruct $\mu_{\theta}(\tilde{x}_{t},t)$, $\forall t\in[0,T]$. We adopt the Mean Squared Error (MSE) as the loss function and introduce regularization terms to prevent overfitting :
@@
Based on the provided calibration data, We can cache the noise prediction outputs of the full-precision model and the quantized model, and then compute $K\in\mathbb{R}^{T\times C}$ offline, which is directly used for quantization errors reconstruction during inference. Different values of $\lambda_{1}$ correspond to different values of $K$. We offer two methods, grid search and empirical rule, to determine the optimal value and compare their advantages and disadvantages. According to the experiments in Appendix F, the latter is adopted, that means $\lambda_{1}=0.01\times\frac{\text{mean}(\tilde{\mu}^{2})}{\text{var}(\mu)}$. Substituting Eq. 14 into Eq. 13, we obtain the final form of the closed-form solution for cumulative error:
@@ 4 Experiments
Table 1: Quantization Performance Comparison of Different Models. SDXL and SDXL-Turbo generate at $512^{2}$ resolution, while PixArt achieves $1024^{2}$. Evaluation metrics include FID (distribution distance) (Parmar et al., 2024), IR (human preference) (Xu et al., 2023), LPIPS (perceptual similarity) (Zhang et al., 2018), and PSNR (numerical fidelity against 16-bit references) (Li et al., 2024b).
[…]
@@ 4.1 Implementation Details
Quantization Scheme. SVDQuant (Li et al., 2024b) introduces an additional low-rank branch that can mitigate quantization challenges in both weights and activations, establishing itself as a new benchmark for PTQ algorithms. In this study, we build our quantization strategy upon it by integrating cumulative error correction mechanisms. This approach ensures that performance comparisons remain unaffected by operator-level quantization configurations. In the 8-bit configuration, our approach employs per-token dynamic quantization for activations and per-channel weight quantization, complemented by a low-rank auxiliary branch with a rank of 16. For the 4-bit configuration, we apply per-group symmetric quantization to both activations and weights, using a low-rank branch with rank 32 and setting the group size to 64. All nonlinear activation and normalization layers are not quantized, meanwhile the inputs of linear layers in adaptive normalization are kept in 16 bits. To comprehensively evaluate the effectiveness of TCEC, we conducted comparative experiments with recent SOTA quantization algorithms across diverse generation tasks.
Image Generation Evaluation. We benchmark TCEC using SDXL (Podell et al., 2023), SDXL-Turbo (Podell et al., 2023) and PixArt models (Chen et al., 2023; 2024) including both the UNet and DiT backbones. SDXL-Turbo uses the default configuration of 4 inference steps, while SDXL employs the DDIM sampler with 50 steps. Since PixArt utilizes the DPM++ solver, we adapted TCEC to this advanced high-order solver to demonstrate its compatibility across different solvers. The detailed validation process is presented in Appendix D. To precompute the channel-wise scaling matrices $K$ and $\lambda_{1}$ mentioned in section 3.3, we randomly sampled 1,024 prompts from COCO dataset (Chen et al., 2015) as the calibration dataset. To evaluate the generalization capability of TCEC, we sample 5K prompts from the MJHQ-30K (Li et al., 2024a) and the summarized Densely Captioned Images(sDCI) (Urbanek et al., 2024) for benchmarking.
Video Generation Evaluation. We apply TCEC to OpenSORA (Zheng et al., 2024), the videos are generated with 100-steps DDIM with CFG scale of 4.0. We evaluate the quantized model on VBench (Huang et al., 2024b) to provide comprehensive results. Following prior research (Ren et al., 2024), we evaluate video quality from three distinct dimensions using eight carefully selected metrics. Aesthetic Quality and Imaging Quality focus on assessing the quality of individual frames, independent of temporal factors. Subject Consistency, Background Consistency, Motion Smoothness, and Dynamic Degree measure cross-frame temporal coherence and dynamics. Finally, Scene Consistency and Overall Consistency gauge the alignment of the video with the user-provided text prompt. We collected 128 samples from WebVid (Nan et al., 2024) as the calibration dataset to calculate $K$ and $\lambda_{1}$.
@@ 4.2 Main Results
Image Generation Evaluation. As shown in Table 1, we conduct extensive experiments at two quantization precisions: W8A8 and W4A4. We observe that, across all precision levels, TCEC achieves better image fidelity and diversity, and it even matches the 16-bit results under W8A8 quantization. For UNet-based models, on SDXL, our W4A4 model substantially outperforms SVDQuant W4A4, the current SOTA 4-bit approach, achieving an sDCI PSNR of 21.9. This even surpasses TensorRT’s W8A8 result of 21.7, demonstrating robust performance under lower-bit quantization. On SDXL-Turbo, MixDQ W4A4 exhibits abnormal FID and IR metrics, indicating quantization failure. This highlights the greater difficulty of quantizing models with a small number of inference steps. Our W4A4 model surpasses SVDQuant by 0.017 and 0.02 in MJHQ IR and sDCI IR metrics, respectively, suggesting a stronger alignment with human preferences. The larger performance gains of TCEC on SDXL than SDXL-Turbo highlight the critical role of cumulative error correction in longer inference sequences. For DiT-based model, on PixArt-$\sum$, our W4A4 model significantly surpasses SVDQuant’s W4A4 results across all metrics. Leveraging the DPM++ solver, PixArt-$\sum$ demonstrates TCEC’s robustness across different solver configurations. As shown in Figure 2, when compared with ViDiT-Q W4A4 and SVDQuant W4A4, the TECE W4A4 method demonstrates less quality degradation and smaller changes in image content. More visual results can be found in Appendix H.
@@
Video Generation Evaluation. As shown in Table 2, our evaluation compares SOTA PTQ methods under both W8A8 and W4A8 configurations. In Imaging Quality, TCEC achieves 65.56 (W8A8) and 64.97 (W4A8), surpassing ViDiT-Q by +2.08 and +3.90 absolute points respectively. It demonstrates significant advantages in frame-wise quality metrics that evaluate static visual fidelity independent of temporal factors. TCEC achieves 96.27 motion smoothness at 8-bit and 96.01 at 4-bit quantization, outperforming ViDiT-Q by 0.13 and 0.32 respectively. This validates the effectiveness of our temporal-channel decoupled compensation strategy in handling error accumulation across denoising steps. The scene consistency metric reaches 39.58 (8-bit) and 37.20 (4-bit), establishing 1.36 and 1.01 improvements over baselines, which confirms stable long-sequence generation through temporal error propagation modeling.
@@
Hardware Resource Savings. TCEC constructs its quantization strategy based on SVDQuant by integrating cumulative error correction mechanisms. As TCEC’s research focus is not on engineering performance, we fully reuse its inference engine, Nunchaku (Li et al., 2024b). As shown in Table 3, experimental tests indicate that the resulting end-to-end (E2E) latency degradation is less than $0.5\%$. Further analysis can be found in the Appendix G. Therefore, on one 12B model, TCEC-W4A4 reduces the memory footprint by $3.5\times$ compared to the FP16 model and accelerates inference by $3\times$ versus NF4 weight-only quantization on the laptop-level RTX 4090, with engineering performance comparable to SVDQuant and significantly outperforming other PTQ algorithms.
@@
Method Bit-width W/A Imaging Quality Aesthetic Quality Motion Smooth. Dynamic Degree BG. Consist. Subject Consist. Scene Consist. Overall Consist. - 16/16 63.68 57.12 96.28 56.94 96.13 90.28 39.61 26.21 Q-Diffusion 8/8 60.38 55.15 94.44 68.05 94.17 87.74 36.62 25.66 Q-DiT 8/8 60.35 55.80 93.64 68.05 94.70 86.94 32.34 26.09 PTQ4DiT 8/8 56.88 55.53 95.89 63.88 96.02 91.26 34.52 25.32 SmoothQuant 8/8 62.22 55.90 95.96 68.05 94.17 87.71 36.66 25.66 Quarot 8/8 60.14 53.21 94.98 66.21 95.03 85.35 35.65 25.43 ViDiT-Q 8/8 63.48 56.95 96.14 61.11 95.84 90.24 38.22 26.06 ViDiT-Q + TCEC 8/8 65.56 57.12 96.27 61.09 96.23 91.34 39.58 26.20 Q-DiT 4/8 23.30 29.61 97.89 4.166 97.02 91.51 0.00 4.985 PTQ4DiT 4/8 37.97 31.15 92.56 9.722 98.18 93.59 3.561 11.46 SmoothQuant 4/8 46.98 44.38 94.59 21.67 94.36 82.79 26.41 18.25 Quarot 4/8 44.25 43.78 92.57 66.21 94.25 84.55 28.43 18.43 ViDiT-Q 4/8 61.07 55.37 95.69 58.33 95.23 88.72 36.19 25.94 ViDiT-Q + TCEC 4/8 64.97 56.90 96.01 59.42 97.01 90.05 37.20 26.20
Table 3: A comparison of inference overhead between SVDQuant and SVDQuant with TCEC.
Model FP16 SVDQuant W4A4 SVDQuant+TCEC W4A4
Flux-dev 1.0 (T = 30) 26.14 s 9.947 s 9.996 s